Assume that the risk-free rate of interest is 5% and the expected rate of return on the market is 12%. A share of stock sells for $59 today. It will pay a dividend of $2 per share at the end of the year. Its beta is 1.3. What do investors expect the stock to sell for at the end of the year?

expected stock price?
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

=5+1.3*(12-5)

=14.1

Required return=(D1/Current price)+Growth rate

0.141=(2/59)+Growth rate

Growth rate=0.141-(2/59)

=0.107101695

Hence expected price=Current price*(1+Growth rate)

=59*(1+0.107101695)

=$65.319